# Bi-directional Deformation for Parameterization of Neural Implicit Surfaces

### v1

@@ A.1 Implementation details
In our framework, the bi-deformation networks $F_{\text{def}}$ and $F_{\text{inv-def}}$, the SDF network $F_{\text{sdf}}$, the texture network $F_{\text{mat}}$, and the appearance network $F_{\text{shd}}$ are all multilayer perceptrons, consisting of 8, 8, 8, 4, and 4 layers, respectively. Each of these hidden layers has 256 neurons. The latent shape codes $z_{s}$ and the appearance codes $z_{a}$ are 128-dimensional. Within the neural rendering process, we set the frequency of position encoding for positions and viewpoints as $6$ and $4$, respectively. In each iteration, we randomly sample 1024 pixels from each input image for training. As for other parameter settings related to neural field reconstruction, we refer mainly to the work of VolSDF  (Yariv et al., 2021).
Baseline implementation details
@@
For static settings, we remove the warp module in the NeP (Ma et al., 2022). Then we train NeP under two settings, one is under UV ground truth as supervision and the other is not( See Figure 11). The parameter configurations for other settings can be found in the referenced paper.
A.2 Additional results
Our method is capable of supporting parameterization between more general man-made models and polycubes, as shown in Figure  10(a). We have tested this on the DTU dataset (Jensen et al., 2014). Moreover, our parameterization framework can be used to seamlessly stitch segmented texture maps together, as shown in Figure 10(b). In addition to texture and shading, our framework also can edit geometric details on original models, as illustrated in Figure 10(c).
Refer to captionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption (a)
Refer to captionRefer to captionRefer to captionRefer to caption (b)
Refer to caption (c)
Figure 10: Additional results: (a) normals, rendering and parametric domain on DTU dataset (Jensen et al., 2014) (b) Stitching texture. (c) Geometric detail editing.
@@ A.2 Additional results
Method Repre Prior Rendering Regularizer Texture Domain
NeuTex (Xiang et al., 2021) Density init UV-inv V.R. NA Entangled Sphere
ISO-UV (Sagnik Das & Samaras, 2022) SDF init UV D.R. Jacobian Entangled $\mathbb{R}^{2}$
NeP (Ma et al., 2022) density Mesh&UV V.R. Angle Disentangled $\mathbb{R}^{2}$
@@
A.3 Qualitative Comparison
We summarize some parametrization methods(refer to Table 1) to illustrate the effectiveness and superiority of our approach. As shown in Table 2, we categorize these methods into three groups: scene-level, object-level, and pixel-level editing. Scene-level editing methods focus on the entire appearance of a scene including lighting and material. Object-level editing methods employ distinct part-based latent codes to manipulate different attributes of a scene, such as hairstyles in i3DMM (Yenamandra et al., 2021). Pixel-level editing, on the other hand, offers a fine-grained editing result by taking into account precise user guidance.
Method Level Retraining 3D input
EditNeRF (Liu et al., 2021) Object Y N
StylizedNeRF (Huang et al., 2022) Scene N N
i3DMM (Yenamandra et al., 2021) Object N Y
NeuMesh (Yang et al., 2022) Pixel Y N
Seal3D (Wang et al., 2023) Pixel Y N
NeuTex (Xiang et al., 2021) Pixel N Y
NeP (Ma et al., 2022) Pixel N Y
Ours Pixel+Object N N

### v2

@@ Appendix A Implementation Details
In our framework, the bi-deformation networks $F_{\text{def}}$ and $F_{\text{inv-def}}$, the SDF network $F_{\text{sdf}}$, the texture network $F_{\text{mat}}$, and the appearance network $F_{\text{shd}}$ are all multilayer perceptrons, consisting of 8, 8, 8, 4, and 4 layers, respectively. Each of these hidden layers has 256 neurons. The latent shape codes $z_{s}$ and the appearance codes $z_{a}$ are 128-dimensional. Within the neural rendering process, we set the frequency of position encoding for positions and viewpoints as $6$ and $4$, respectively. In each iteration, we randomly sample 1024 pixels from each input image for training. As for other parameter settings related to neural field reconstruction, we refer mainly to the work of VolSDF  [61].
Baseline implementation details.
@@
For static settings, we remove the warp module in the NeP [27]. Then we train NeP under two settings, one is under UV ground truth as supervision and the other is not (see Figure 13). The parameter configurations for other settings can be found in  [27].
@@
FS-Syn dataset preparation.
We derived a customized dataset from Facescape [58], as shown in Figure 14. We randomly chose 10 human heads with neutral expressions from uniformly topologized models in the FaceScape dataset. Subsequently, we illuminated the human head model by moving both the light source and the camera, generating diverse lighting effects. Following this, we captured 30 images from different angles around the human head, each accompanied by its corresponding camera parameters.
Figure 14: We create the FS-Syn dataset from the FaceScape dataset. We illuminated the human head by moving both the light source and the camera, creating varying lighting effects.
Appendix B Additional Results
We also tested our method on the DTU dataset [17] as shown in Figure  15(a). Moreover, our parameterization framework can be used to seamlessly stitch segmented texture maps together, as shown in Figure 15(b). In addition to texture and shading, our framework also can edit geometric details on original models based on parameterization results and the assigned displacement map, as illustrated in Figure 15(c).
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
(c) Refer to caption Refer to caption Refer to caption
Figure 15: Additional results: (a) Visualization of normal maps, neural rendering results, and parametric domain on models from the DTU dataset [17]. (b) Textures. The textures associated with the original meshes (middle) are discontinuous and exhibit significant distortion, making them less suitable for editing purposes. In contrast, the textures resulting from our parameterization (right) are globally continuous and exhibit minimal distortions. (c) Geometric detail editing. (d) Source geometry, target appearance, and texture transfer results. Durian and Apple undergo joint training as they belong to the same category.
Appendix C Geometry Editing via Modifying the Parametric Domain $\mathcal{D}$
Figure 16: Pipeline of geometry editing on the parametric domain. We replace the original parametric domain with a modified one and immediately render the edited objects.
In addition to the 2D displacement map, we also provide a geometry-editing method (See Figure 16) based on the parametric domain. As shown in Figure 17, we remove the handle of the cup in the parametric domain, and then we render the cup model in volume rendering without any training again.
Figure 17: Editing the geometry of the parametric domain $\mathcal{D}$ naturally induces a modified neural implicit surface. We edit the geometry of the cup by removing and deforming the handle in $\mathcal{D}$. The top row shows the original parametric domain $\mathcal{D}$ and two modified domains. The bottom row shows the neural implicit surface $\mathcal{S}$ corresponded to each parametric domain.
Method Level Retraining 3D input
EditNeRF [25] Object Y N
StylizedNeRF [16] Scene N N
i3DMM [63] Object N Y
NeuMesh [57] Pixel Y N
Seal3D [48] Pixel Y N
NeuTex [53] Pixel N Y
NeP [27] Pixel N Y
Ours Pixel+Object N N
Appendix D Qualitative Comparison
We summarize some parametrization methods (refer to Table 2) and some neural editing methods (refer to Table 3) to illustrate the effectiveness and superiority of our approach. As shown in Table 1, we categorize these texture editing methods with other editing methods into three groups: scene-level, object-level, and pixel-level editing. Scene-level editing methods focus on the entire appearance of a scene including lighting and material. Object-level editing methods employ distinct part-based latent codes to manipulate different attributes of a scene, such as hairstyles in i3DMM [63]. Pixel-level editing, on the other hand, offers a fine-grained editing result by taking into account precise user guidance.
Method Representation Prior Rendering Regularization Texture Domain
NeuTex [53] Density Init. UV-inv V.R. NA Entangled Sphere
ISO-UV [39] SDF Init. UV D.R. Jacobian Entangled $\mathbb{R}^{2}$
NeP [27] Density Mesh & UV V.R. Angle Disentangled $\mathbb{R}^{2}$
@@ Appendix D Qualitative Comparison
Method Proxy Deformation Removal Texture Retraining Category
NeRF-Editing [64, 65] Mesh Y N N N N
NeuMesh [57] Mesh Y N Y Y N
DRFC [55] Cage Y N N N N
Ours Parametric domain Y Y Y N Y
Table 3: Qualitative comparison with other methods for editing the geometry of neural radiance fields. While all these methods exhibit deformation capabilities, they encounter challenges of removing specific parts, such as a cup’s handle, as illustrated in Figure 17. By editing the parametric domain $\mathcal{D}$, we can immediately obtain the deformed neural implicit surfaces through inference on the modified domain $\mathcal{D}^{\prime}$. Our method is trained on a specific category of objects, such as human heads and fruits. This specialized training enables our approach to transfer textures and geometries among objects in the same category.
GJBLBZDJBs9mE4zjwfZ85lAGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">